the prairie land base ball conference has 6 teams the teams name are listed below.
*Aces
*Broncos
*Cougars
*Wildcats
*Eagles
*Mustangs
How many games must be played so that every team plays each other once?

Answer by stanbon(75887)

the prairie land base ball conference has 6 teams the teams name are listed below.
*Aces
*Broncos
*Cougars
*Wildcats
*Eagles
*Mustangs
How many games must be played so that every team plays each other once?
----
2 teams: 1 game = 2C2
3 teams: ab,ac,bc : 2+1 = 3 games = 3C2
4 teams: ab,ac,ad,bc,bd,cd = 3+2+1 = 6 games = 4C2
5 teams: 4+3+2+1 = 10 games = 5C2
...
6 teams: 5+4+3+2+1 = 15 games = 6C2
=======================================
Cheers,
Stan H.